Let K be a field and (c,d) a pair of relatively prime positive integers. We
 explain the computation of the Grothendieck group of algebras of the type
K(z,x_1,x_2,...x_n), with relation z power c=f(x_1,x_2...x_n) where f is isobaric
                               of weight d.